OpenStudy (anonymous):

Integrate ((9sin^4(t) + 9cos^4(t))^1/2

OpenStudy (anonymous):

It's all underneath the square root

OpenStudy (luigi0210):

This right? \(\Large \int (\sqrt{9sin^4 t +9cos^4 t}) dt\)

OpenStudy (anonymous):

Yes

OpenStudy (luigi0210):

I would trying reducing it down, factor out that 9 first: \(\Large \int \sqrt{9(sin^4 t +cos^4 t}) dt \) Since 9 can come out of the square roots, you get this: \(\Large \int 3(\sqrt{sin^4 t +cos^4 t}) dt \) Make sense so far?

OpenStudy (anonymous):

Yep got that down already - then I tried to use half angle identities but got stuck there

OpenStudy (luigi0210):

You factored a square from each right?

OpenStudy (luigi0210):

\(\Large 3\int (\sqrt{(sin^2 t )^2+(cos^2 t)^2} dt \) ?

OpenStudy (anonymous):

Yep then used 1+cos(2t)/2 and 1-cos(2t)/2

OpenStudy (anonymous):

???

OpenStudy (luigi0210):

But yea, what did you get after you tried that?

OpenStudy (anonymous):

Tried expanding them - but since its a square under square root can be factored out?

OpenStudy (anonymous):

??

OpenStudy (emmatassone):

i dont want to be killjoy but i think that integral dont have solution. its similar to:\[\int\limits_{}^{}\sqrt{a.\cos^{2}(x)+b.\sin ^{2}(x)} And this integral dont have solution neither.Its called the elliptic integral.

OpenStudy (luigi0210):

Yea that's what I was getting too

OpenStudy (luigi0210):

And your num lock is off or something there? :P

OpenStudy (anonymous):

I am doing arc length it's from 0 to pi/2

OpenStudy (luigi0210):

That would of been a bit helpful to know

OpenStudy (anonymous):

Answer in back says 3/2

OpenStudy (emmatassone):

Maybe its a numerical solution or another way to calculate it, because you cant calculate it direclty with integrating methods.

OpenStudy (irishboy123):

you should post or scan the question it looks like an arc lngth, but why guess? :p

OpenStudy (amistre64):

might help to post the actual question ... or scan yeah

OpenStudy (anonymous):

Ok will try to write it out am on my iPad so here it goes Find the length of the following r(t) = <cos^3(t), sin^3(t)

OpenStudy (anonymous):

So found derivative and then did magnitude and got to where the question is now !! :)

OpenStudy (anonymous):

:-/

OpenStudy (amistre64):

what techniques have you covered?

OpenStudy (anonymous):

I had thought to use the half angle formula to then take integral of ! In class covered take r'(t) then magnitude of r'(t) then take integral

OpenStudy (emmatassone):

i think you have derivated wrong, check it again

OpenStudy (anonymous):

Cos^3(t) = -3sin^2(t) Sin^3(t) = 3cos^2(t)

OpenStudy (amistre64):

im thinking .. and it may not be any better y = sin^3(acos(cbrt(x))) to eliminate the parameter maybe?

OpenStudy (emmatassone):

you are forgeting chain rule

OpenStudy (amistre64):

accursed chain rule lol

OpenStudy (amistre64):

writing it as say: (cos(t))^3 might make it easier to recall chaining it

OpenStudy (anonymous):

Then what is it !? There is nothing with the t to do chain rule

OpenStudy (amistre64):

if u = cos(t) u^3 derives to 3u^2 u'

OpenStudy (emmatassone):

\[(\cos ^{3}(x))\prime=3.\cos ^{2}(x).(-\sin(x))\]

OpenStudy (emmatassone):

same with the sine

OpenStudy (anonymous):

So sine = 3sin^2(t)cos(t)

OpenStudy (anonymous):

Gosh dang chain rule lol

OpenStudy (emmatassone):

yep, it was a calculation mistake lol

OpenStudy (anonymous):

Wow lol such stupid mistake lol